Please help solve: The mean weight of 9 pumpkins is 16 pounds. One of the pumpkins is removed. The mean weight of the remaining 8 pumpkins is 13 pounds. What is the weight of the pumpkin that was removed, in pounds?
Answer by asha(30)

The mean weight of 9 pumpkins is 16 pounds.
the total weight of 9 punpkins is 16*9 =144 pounds.
The mean weight of 8 pumpkins is 13 pounds.
the total weight of 8 pumpkins is 13*8 =104 pounds.
The weight of the pumpkin that was removed is
144-104=40
The answer is 40 pounds.